OpenStudy (anonymous):

please help me out. :) what is the polar form for y = 3x ?

ganeshie8 (ganeshie8):

\(x = r \cos \theta\) \(y = r \sin \theta\)

ganeshie8 (ganeshie8):

substitute these in the given equation

OpenStudy (anonymous):

r sin theta = 3 r cos theta ?

ganeshie8 (ganeshie8):

yup ! u may cancel r both sides

ganeshie8 (ganeshie8):

\(\sin \theta = 3 \cos \theta\) \(\tan \theta = 3\)

OpenStudy (anonymous):

sin theta = 3 cos theta ?

ganeshie8 (ganeshie8):

yes, divide cos theta both sides

OpenStudy (anonymous):

tan theta = 3 ?

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

but why? i thought in the polar form, there's always "r" in the final answer..

ganeshie8 (ganeshie8):

in cartesian form, do u have \(y\) in the equation always ?

OpenStudy (anonymous):

am, i don't know.. maybe no?